OpenStudy (anonymous):

Find the equation of the quadratic function with zeros 10 and 14 and vertex at (12, -8). y = 2x2 + 48x + 280 y = -2x2 - 48x + 280 y = -2x2 + 48x - 280 y = 2x2 - 48x + 280

OpenStudy (zale101):

i factored it but i didn't got one of your choices for some reason (x-10)(x-14) I think after you factor it u substitute the vertex at (11 , -2) y=a((x-10)(x-14)) but that's not the end though

OpenStudy (zale101):

but make sure u factor (x-10)(x-14) first

OpenStudy (zale101):

and then substitute

OpenStudy (anonymous):

I'm not sure how to do that. I am unfortunately pretty bad at math :/

OpenStudy (zale101):

factor (x-10)(x-14), can you factor it out by using the foil method?

OpenStudy (anonymous):

2x-14x-10x+140?

OpenStudy (zale101):

Well, FOIL stands for First, Outer, Inner, Last. (x-10)(x-14) =(x · x+x · -10-14 · x-14 · -10)

OpenStudy (zale101):

2x-14x-10x+140 is wrong, u first multiply the first factors